Determine death rate for USA for guns. 100,000 people rate

America:

[32,351 deaths]

[311.7 million people]

this equates to 3117 groups of 100,000

What is the rate per 100,000 people to die from guns?

Let’s compare the US to Japan. The firearm death rate of Japan was what percent of the US rate? Round your answer to 2 decimal places.

Japan had the lowest rate, at

0.05 gun deaths

per 100,000
Answer #1

Gun deaths per 100,000 in the US = 32351 / 3117 = 10.37899 = 10.38 per 100,000

what % of US death rates was the death rate for Japan

Therefore (% / 100) * Total = A Portion, where Total is the number always followed by of.

Therefore (% / 100) * 10.38 = 0.05

Therefore % = (0.05 100) / 10.38 = 0.48%